ABC Company sells 19,996 chairs a year at an average price per chair of $13.26. The carrying cost per unit is $9.88. The company orders 108 chairs at a time and has a fixed order cost of $68.78 per order. The chairs are sold out before they are restocked. What are the total costs if the company orders the optimal quantity (Economic Order Quantity)?

Enter your answer rounded off to two decimal points. Do not enter comma or $ in the answer box.
Answer #1

2 * Total Demand * Costper Order EconomicOrder Quantity = 1 Holding Cost

EconomicOrder Quantity = 1 2 * 19996 * 68.78 9.88

= 527.61 OR 258 Chair

No of Orders = Demand / EOQ

= 19996/ 258

= 77.50 OR 78 Orders

Average Inventory = EOQ/2

= 258/2

= 129 Chairs

Order Cost now = 78 orders * 68.78 = 5364.84

Carrying Cost = Avg Inventory * Carrying Cost

= 129 Chairs * 9.88 = 1274.52

Then Total Cost will be = 5364.84 + 1274.52 = 6639.36

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.

2 * Total Demand * Costper Order EconomicOrder Quantity = 1 Holding Cost

EconomicOrder Quantity = 1 2 * 19996 * 68.78 9.88